### v1

# Chelsea Zou    Kenneth J. Kurtz

Abstract
We introduce a cluster-based generative image segmentation framework to encode higher-level representations of visual concepts based on one-shot learning inspired by the Omniglot Challenge. The inferred parameters of each component of a Gaussian Mixture Model (GMM) represent a distinct topological subpart of a visual concept. Sampling new data from these parameters generates augmented subparts to build a more robust prototype for each concept, i.e., the Abstracted Gaussian Prototype (AGP). This framework addresses one-shot classification tasks using a cognitively-inspired similarity metric and addresses one-shot generative tasks through a novel AGP-VAE pipeline employing variational autoencoders (VAEs) to generate new class variants. Results from human judges reveal that the generative pipeline produces novel examples and classes of visual concepts that are broadly indistinguishable from those made by humans. The proposed framework leads to impressive but not state-of-the-art classification accuracy; thus, the contribution is two-fold: 1) the system is uniquely low in theoretical and computational complexity and operates in a completely standalone manner compared while existing approaches draw heavily on pre-training or knowledge engineering; and 2) in contrast with competing neural network models, the AGP approach addresses the importance of breadth of task capability emphasized in the Omniglot challenge (i.e., successful performance on generative tasks). These two points are critical as we advance toward an understanding of how learning/reasoning systems can produce viable, robust, and flexible concepts based on literally nothing more than a single example.
One-Shot Learning, Omniglot Challenge, Gaussian Mixture Models, Variational Autoencoders, Concept Learning
1 Introduction
The ability of humans to acquire novel concepts after minimal exposure to examples is an important constituent of general intelligence. Humans have the remarkable ability to quickly abstract concepts and extrapolate from one or few provided examples (Lake et al., 2015), thereby allowing for efficient and adaptable learning and reasoning. On the contrary, most current machine learning (ML) architectures require large amounts of data to learn, massive numbers of parameters (e.g., GPT-3: 175B, AlexNet: 62.3M, VGG16: 138M), and in many cases pre-training or access to external data or trained models (Hendrycks et al., 2019; Han et al., 2021; L’heureux et al., 2017; Zhou, 2016). Hence, a key computational challenge is to understand how an intelligent system with minimal complexity and without reliance on external supports can acquire new concepts from highly restricted training data (Chollet, 2019).
In this paper, we approach the challenges of one-shot learning (Wang et al., 2020; Kadam & Vaidya, 2020) by developing a framework that can perform the classification and generative tasks defined by the Omniglot challenge of handwritten characters—a testbed designed to promote the study of human-like intelligence in artificial systems (Lake et al., 2015, 2019). In the classification task, a single image of a novel character is presented and the aim is to correctly identify another instance of that character from a choice set of characters. In the generative tasks, the goal is to create new variants of characters that are indistinguishable from human drawings. While the classification task has received much attention, there has been limited success in the attempt to achieve both types of tasks with the same model (despite an emphasis on exactly this breadth of functionality in the Omniglot challenge).
The Omniglot challenge is a highlight of the emerging field of computational cognition. Unlike traditional machine learning research, the goal is not just to achieve high classification performance on novel test items given a set of labeled training data. The Omniglot challenge differs in ways that derive from perceived weakness in dominant ML approaches relative to human intelligence: 1) the ability to form a concept from a single training item rather than expansive training sets; and 2) the ability to form flexible, robust, and transparent/trustable concepts. Accordingly, the challenge focuses on one-shot learning under a broad, multi-task interpretation of inductive concept learning that asks proposed systems not only to classify new items, but also to successfully perform a set of generative tasks. The latter includes inventing viable new instances of an individual character in an alphabet, inventing viable new characters within an alphabet, and inventing viable new characters without constraint (note: an additional proposed task involves parsing characters into ordered strokes, however we see this as less central to the force of the challenge). An important intellectual thread embedded in the Omniglot challenge is a question at the heart of computational cognition: Are structured (also referred to as compositional, symbolic, or causal) representations necessary to learn and reason with concepts in a sophisticated manner? We attempt to avoid a black-and-white approach to this question. Rather than committing to massive data and architectures to estimate models that yield high performance in tightly constrained evaluations or assuming the availability of powerful symbolic systems prior to new learning, we investigate design principles to construct quasi-structured representations of a concept built from nothing more than a single training item.
Figure 1: The raw image is shown on the left. The inferred clusters of the GMM are shown in the middle. The newly generated abstracted Gaussian prototype (AGP) is sampled from the inferred parameters.
To address both classification and generative tasks of the Omniglot challenge, we propose the Abstracted Gaussian Prototype (AGP) which leverages Gaussian Mixture Models (GMMs) to flexibly model visual concepts of handwritten characters. This is accomplished by capturing character subparts as inferred Gaussian components that flexibly represent ’what should be where’ for a concept based on a single example. GMMs are unsupervised clustering algorithms that represent data as a finite sum of Gaussian distributions (Duda et al., 1973; Yu et al., 2015; Liang et al., 2022a; Reynolds et al., 2009). A key insight driving our approach is viewing the pixels of an image as instances in the domain of an individual character. We use GMM-based clustering to model the topological subparts of each class in terms of its unique distribution. Based on the parameters underlying each component’s distribution, the GMMs are then used to generate additional model-consistent pixels that augment the subpart representations. Finally, the collective ensemble of these generated subparts form what we refer to as the AGP, see Figure 1.
AGPs provide a way to extrapolate beyond the constraints of a single available instance based on: 1) the clustering functionality that imparts a quasi-structural analysis of the raw input into underlying parts with relative locations; and 2) the generative functionality that enhances the item encoding by extrapolating its underlying model. This is a promising intermediate strategy between the problem of completely lacking a structural representation and that of being overly structured in the sense of requiring the overhead of a built-in representational vocabulary and being tied to fixed expectations. In general, AGPs offer a method for generative image segmentation. The resulting higher-level representation (AGP) is successfully induced from a single example and supports a range of conceptual tasks. AGPs are characterized by two elements: 1) probabilistic capture of data substructures, and 2) parameterized generative modeling. The first element allows the system to accurately model diverse concepts and robustly handle noise. The second element allows the creation of sufficiently flexible representations of each concept.
To achieve one-shot classification, we leverage the generalized class representation of AGPs in combination with a cognitively-inspired metric to assess the similarity between AGPs. A classic and highly influential psychological theory of human similarity judgement is a set-theoretic approach known as the contrast model (Tversky, 1977). This similarity metric employs an asymmetric weighting of the number of common and distinctive features between being compared. In the context of the one-shot classification task, the features to be evaluated are pixel intersections between AGP representations of the characters (further details below), so a classification decision from a choice set is made based on the highest similarity score to the target item. For the challenging generative tasks in the Omniglot challenge requiring the invention of new character types or tokens, we build on the approach outlined above to devise an AGP-VAE pipeline to create new classes. First, we synthetically generate a diverse training set of AGPs from each provided character. Then, we employ a variational autoencoder (VAE) (Kingma & Welling, 2013), a type of generative neural network architecture, to learn a continuous latent space that encapsulates a probabilistic distribution over different classes of generated AGP train sets. This novel AGP-VAE pipeline has the ability to interpolate between subparts of the discrete prototypes of the AGPs by sampling from a global feature space that encapsulates the local features of different classes. Instead of relying on the independent models of each class, we are able to represent the available classes in a subspace that can be sampled to produce novel character classes that conform to the distributional characteristics of provided data. In sum, the key design principles of our contribution are as follows:
  1. 1.
    We formulate the framework of an AGP as a generative image segmentation method that provides a higher-level representation from a single character instance by applying a clustering approach at the pixel level.
  2. 2.
    We make novel use of a cognitively inspired similarity metric between AGP representations to perform one-shot classification.
  3. 3.
    We employ a novel AGP-VAE pipeline for generative tasks to create new visual concepts.
Our framework provides a flexible and robust way to handle both classification and generative tasks. The system learns and performs these tasks without reliance on pre-training or a pre-formed symbolic system. Further, the approach is highly transparent and relies on a small number of clear design principles that are novel applications of established computational constructs.
2 Prototypes in Human Concept Learning
Concepts are mental representations acquired through inductive learning processes that are used to categorize and reason about objects, events, and relational situations (Gregory, 2002). Psychologists have emphasized that concepts are formed and represented through the integration and combination of simpler parts known as features or attributes (Schyns et al., 1998; Farhadi et al., 2009). In the concepts and categories literature in cognitive psychology, a prominent account is prototype theory in which the mental representations of categories consist of a stored summary of the central tendency of feature values across observed members of a category (Rosch, 1973; Hampton, 2006; Posner & Keele, 1968; Minda & Smith, 2011). The prototype of a concept is a statistical average or an abstraction of all instances observed as members of a category. Categorization then operates as a matter of finding the best match among candidate prototypes which supports graded membership and classification decisions based on family resemblance rather than logical rules (Rosch & Lloyd, 1978; Mervis & Rosch, 1981). This approach has contrasted sharply in the psychological literature with exemplar-based similarity approaches that eschew explicit abstraction in favor of storing labeled instances (Nosofsky, 1988, 1986). The main advantage of prototype theory is its broad computational and psychological plausibility: prototypes are economical in terms of representational and processing requirements, and they capture the intuition that the generic meaning underlying a category is represented explicitly and independently of its members.
In our framework, we adapt a form of prototype theory to address the Omniglot challenge. Prototypes are not necessarily formed by storing the centroid of category members but can instead be realized implicitly through statistical learning or neural network architectures like simple linear associators or auto-associative systems which adapt their weights to capture the statistical regularities of a domain (Biehl et al., 2016; Rosenblatt, 1958; Ruck et al., 1990). To be clear, prototypes are nearly always a matter of abstracting a summary representation or model from individual cases. By contrast, our approach to one-shot learning transforms a single example into a prototype by taking a particular configuration of pixels as the basis for a set of probabilistic clusters that imply an underlying distribution with a generative capacity. The crux of our approach lies in the formation of an ensemble of augmented subparts achieved by using the parameters inferred from the Gaussian components of the GMMs. The abstracted Gaussian prototype is an abstraction from a single case in order to create its own subparts via clustering, establish the distributional central tendency and variability of its subparts, inherently capture spatial relational properties between the subparts, and further populate the prototype with distribution-consistent generated pixels. The result of this process can be seen as similar to a summary representation of the central tendency across varying training examples with the additional quasi-structural aspect of an underlying model that captures the form and location of subparts (without any explicit propositional or symbolic representation).
Just as the prototype theory of human categorization relies on similarity to the stored prototype in order to determine likelihood of category membership, we invoke a psychological similarity metric for classification relative to AGPs. Tversky’s (1977) model of similarity proposed that individuals assess similarity by considering the number of featural differences and commonalities between items along with an additional design principle of highlighting (via greater weighting) the importance of the differences. The Tversky index is given by:
$T(A,B)=\frac{|A\cap B|}{|A\cap B|+\alpha|A\backslash B|+\beta|B\backslash A|}$ (1)
where $|A\cap B|$ is the size of the intersection of sets $A$ and $B$, $|A\backslash B|$ is the size of the set difference of $A$ without $B$, $|B\backslash A|$ is the size of the set difference of $B$ without $A$, and $\alpha,\beta$ are non-negative parameters controlling the weight given to differences in the two sets. We employ a slightly simpler metric that has only one parameter to control the magnitude of the penalty to the set differences in comparing the intersections of AGPs, described in Section 5.1.2.
3 Related Works
3.0.1 One-Shot Classification:
Many neural-based models have been successful at one or few-shot classification tasks (Finn et al., 2017; Santoro et al., 2016; Salakhutdinov et al., 2012). For instance, Siamese Neural Networks involve twin sub-networks sharing the same parameters that are trained to learn embeddings capturing the similarity or dissimilarity between pairs of instances (Koch et al., 2015; Chicco, 2021). Another approach through Prototypical Networks learn a representative prototype for each class based on the mean of the embeddings in the latent space and classify according to these distances (Snell et al., 2017). Similarly, Matching Networks work by employing an attention mechanism on embeddings of the labeled set of instances to forecast classes for unlabeled data (Vinyals et al., 2016). However, all of these neural-based classification approaches require an initial training phase for the network to learn a general understanding of the task. Furthermore, they are incapable of addressing generative tasks. Our proposed approach, on the other hand, offers a direct way for both classification and generative tasks to learn specific concepts from one, and only one, shot without background training on other data.
3.0.2 One-Shot Generation:
One recent approach introduces GenDA for one-shot generative domain adaptation using pre-trained Generative Adversarial Networks (Yang et al., 2023; Goodfellow et al., 2014). GenDA designs an attribute classifier that guides the generator to optimally capture representative attributes from a single target image, in turn, synthesizing high-quality variant images. However, this approach relies on source models that are pre-trained on large-scale datasets such as FFHQ and Artistic-Faces dataset.
3.0.3 Bayesian Models:
Significant progress has been made to address both one-shot classification and generative tasks through Bayesian implementations, such as the Object Category Model (Fei-Fei et al., 2006) involving parametric representations of objects and prior knowledge when faced with minimal training examples. These Bayesian principles are manifested in the approach proposed by the original authors for the Omniglot dataset (Lake et al., 2011). In this system, a stroke model learns part-based representations from previous characters to help infer the sequence of latent strokes in new characters. An extension of this work introduces Bayesian Program Learning (BPL), which learns a dictionary of sub-strokes and probabilistically generates new characters by constructing them compositionally from constituent parts and their spatial relationships (Lake et al., 2015). BPL and the stroke model, however, requires the model to learn from stroke-data trajectories in order to extract and store a dictionary of primitive parses at the sub-stroke level. While this model first requires information from live-drawings, the approach may be unfeasible when temporally labeled sequential stroke data is not accessible. In contrast, our approach uses generative image segmentation to directly infer the sub-strokes of the characters using GMMs which allows our model to learn purely from a single raw image.
In essence, existing approaches rely on either pre-training and/or built-in representational systems that fall outside of the scope of our stricter interpretation of the challenge; and many do not address the breadth of tasks in Omniglot. For instance, (Rezende et al., 2016) involves pre-training as part of its methodology to achieve one-shot generalization, which is a common approach in deep generative models to understand and generalize from very few examples. In (Edwards & Storkey, 2016), the concept of the neural statistician implies learning statistical representations from data, which also involve pre-training to capture the statistical properties across different contexts or datasets. In (Liang et al., 2022b), given its neuro-symbolic approach, this paper uses built-in knowledge/representational languages which involves more explicit and expensive forms of knowledge representation. Finally, in (Giannone & Winther, 2022; Antoniou et al., 2017; Boutin et al., 2022), these papers use highly complex methodologies and low transparency architectures like GANS and hierarchical approaches that attempt to model data at multiple levels of abstraction. While the performance of these works offer solutions to few shot learning, our work attempts to capture a different focus. Our system is trained on nothing except the presented single characters for each task.
4 Background
In this section, we provide the mathematical background underlying GMMs and VAEs.
4.1 Gaussian Mixture Models
A GMM is a probabilistic clustering model that assumes the data is generated from a combination of multiple Gaussian distributions (Duda et al., 1973). Each Gaussian component $k\in K$ represents a cluster in the dataset, and is characterized by its unique parameters mean $\mu$, standard deviation $\sigma$, and a weight $\pi$. A univariate Gaussian probability density function (PDF) for random variable $X$, which represents the probability of observing the given datapoint, is defined as the following:
$P(X|\mu,\sigma)=\mathcal{N}(\mu,\sigma)=\frac{1}{\sigma\sqrt{2\pi}}\exp\left({\frac{(X-\mu)^{2}}{2\sigma^{2}}}\right)$ (2)
Similarly, a multivariate Gaussian PDF is defined as $\mathcal{N}(\mu,\Sigma)$:
$=\frac{1}{(2\pi)^{d/2}|\Sigma|^{1/2}}\exp\left(-\frac{1}{2}(X-\mu)^{T}\Sigma^{-1}(X-\mu)\right)$ (3)
where $X$ is now a d-dimensional feature vector, $\mu$ is a d-dimensional vector representing the means of the distribution, and $\Sigma$ is a d x d covariance matrix. A finite GMM can be expressed by a weighted sum over K components:
$P(X|\pi,\mu,\Sigma)=\sum_{k=1}^{K}\pi_{k}*\mathcal{N}(X|\mu_{k},\Sigma_{k})$ (4)
satisfying the condition where:
$\sum_{k=1}^{K}\pi_{k}=1$ (5)
4.2 Variational Autoencoders
Variational autoencoders (VAEs) are neural networks characterized by an encoder-decoder architecture. Unlike traditional autoencoders (Bank et al., 2023), VAEs are equipped with a probabilistic framework based on variational inference. This framework employs networks to approximate otherwise computationally intractable posterior distributions, enabling the model to learn continuous representations of discrete input classes (Kingma & Welling, 2013; Kingma et al., 2019). This latent space facilitates the generation of diverse and novel samples, making VAEs a versatile tool for tasks such as image generation.
Encoder: The encoder of a VAE maps input data $x$ to a latent space variable $z$, and is defined by an approximate posterior distribution:
$q_{\phi}(z|x)=\mathcal{N}(z;\mu_{z}(x),\sigma_{z}(x)^{2})$ (6)
where $\mu_{z}(x)$ and $\sigma_{z}(x)$ are the mean and standard deviation of the approximate posterior learned by the encoder’s weights and biases $\phi$.
Sampling with Reparametrization Trick: To obtain a sample $z\sim q_{\phi}(z|x)$ from the latent space learned by the encoder, the reparametrization trick is used to maintain differentiability for backpropagation:
$z=\mu_{z}(x)+\sigma_{z}(x)\odot\epsilon$ (7)
where $\epsilon$ is typically sampled from a fixed standard normal distribution:
$\epsilon\sim\mathcal{N}(0,1).$ (8)
Decoder: The decoder of a VAE with parameters $\theta$ maps a sample $z$ from the latent space back to the original feature space, generating a reconstruction $\hat{x}$ from the conditional likelihood distribution:
$p_{\theta}(x|z)=\mathcal{N}(\hat{x};\mu_{x}(z),\sigma_{x}(z)^{2})$ (9)
where $\mu_{x}(z)$ and $\sigma_{x}(z)$ are the mean and standard deviation of the reconstructed data.
Loss Function: The training objective for VAEs is to maximize the Evidence Lower Bound (ELBO). The ELBO is defined by:
$\mathcal{L}(\phi,\theta)=E_{z\sim q_{\phi}(z|x)}[\log p_{\theta}(x|z)]-\text{$D_{KL}$}(q_{\phi}(z|x)||p(z))$ (10)
where the first term is the expected value of the log-likelihood of $x$ given $z$ and the second term is the Kullback-Leibler (KL) divergence between the approximate posterior and the prior distribution. The KL divergence measures the difference between two probability distributions and acts as a latent space regularization term that encourages the learned approximate posterior space to be close to the true posterior. Let $J$ be the dimensionality of $z$. For the case of two Gaussian distributions, it is defined as:
[…]
Here, $p(z)$ $\sim$ $\mathcal{N}(0,1)$, which is a classical assumption in a standard VAE.
5 Approach
In this section, we describe and formalize our approach to the classification and generative tasks in the Omniglot Challenge. The Omniglot dataset consists of 1623 hand-written characters taken from 50 different alphabets, with 20 examples for each class (Lake et al., 2019). All code, question sets, and data can be accessed at https://github.com/bosonphoton/AbstractedGaussianPrototypes
5.1 Classification Task
In a one-shot classification task, there is a set of $N$ classes, denoted as $\mathcal{C}=\{c_{1},c_{2},\ldots,c_{N}\}$, and one available instance of each class. The given set of these single instances is denoted as $\mathcal{X}=\{x_{1},x_{2},\ldots,x_{N}\}$. Let $f:\mathcal{X}\rightarrow\mathcal{C}$ be the function that maps an instance to a corresponding class. Given an unseen instance $q$ (usually referred to as the query) of an existing class, the goal of one-shot classification is to correctly determine the class of the query instance: $c_{q}$ = $f(q)$, where $c_{q}\in\mathcal{C}$. Our classification function is comprised of two steps: (1) AGP generation, and (2) the cognitively inspired similarity metric to determine classification choice. First, a separate GMM is used to model each concept, where each cluster of the model is assumed to represent a unique subpart of the concept. Next, we generate new abstracted subparts by probabilistically sampling new pixels from the inferred clusters. The collective ensemble of these generated subparts form the AGP, a higher level representation for each class, denoted $\mathcal{P}$. Using these prototypes, we use the similarity metric based on Tversky’s (1977) contrast model to compute the match between the query prototype $\mathcal{P}_{q}$ and the prototype of each instance in the starting set $\mathcal{P}_{i}$ for $i\in\mathcal{X}$. The class which produces the highest similarity score is selected to be the class of $q$. The details of each step are formalized below.
5.1.1 Abstracted Gaussian Prototype (AGP) Generation
Each instance of a concept is provided as a binary image of pixels. Under the probabilistic framework of a GMM, let us define each sampled pixel as the realization of a random variable, characterized by its PDF corresponding to the inferred Gaussian component. Each instance in $\mathcal{X}$, along with $q$, is first segmented into its unique component subparts using a GMM, where $\mathcal{G}=\{g_{1},g_{2},\ldots,g_{k}\}$ represents the set of different subparts in each instance and $k$ is a hyperparameter controlling the number of components. Here, $\mathcal{G}$ represents the mixture of Gaussian components of each instance, which allows the GMM to sample from the fitted distribution for each component $g_{i}$ and generate new augmented subparts $p_{i}$. We define the ensemble of these subparts as the prototype $\mathcal{P}$ of the class, where $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{k}\}$.
5.1.2 Similarity Metric
The similarity metric is computed between the query prototype $\mathcal{P}_{q}$ and each prototype of the available instances $\mathcal{P}_{i}$. Each prototype $\mathcal{P}$ is the entire set of its 2D image coordinates, generated by the sampled components in a GMM. To simplify the notation, let $A$ be the set of all image coordinates for $\mathcal{P}_{q}$, and $B$ be the set of all image coordinates for $\mathcal{P}_{i}$. We base our similarity metric on the Tversky index with the following equation:
$S(A,B)=\ |A\cap B|-\beta\ |A\triangle B|$ (12)
where $A\triangle B=(A\backslash B)\cup(B\backslash A)$ is the symmetric difference representing the non-intersections and $\beta>1$ is a weight hyperparameter that ensures a larger penalty of this difference. The 2D image coordinates $a\in A$ and $b\in B$ will be denoted as $(x_{a},y_{a})$ and $(x_{b},y_{b})$, respectively. In order to calculate $S$ by computing the number of intersecting pixels, we consider each pixel as a circle center with radius $r$. This ensures isotropy in our intersection calculation, maintaining consistent measurement of distances between pixels uniformly in every direction. Then, we can define:
$|A\cap B|=\sum_{a\in A}\sum_{b\in B}\begin{cases}1&\text{if }\sqrt{(x_{a}-x_{b})^{2}+(y_{a}-y_{b})^{2}}\leq r\\ 0&\text{otherwise}\end{cases}$ (13)
Using similar notation, it follows that the symmetric difference is:
$|A\triangle B|=\sum_{a\in A}\sum_{b\in B}\begin{cases}1&\text{if }\sqrt{(x_{a}-x_{b})^{2}+(y_{a}-y_{b})^{2}}>r\\ 0&\text{otherwise}\end{cases}$ (14)
Finally, the $\mathcal{P}_{i}$ with the highest similarity score with $\mathcal{P}_{q}$ is deemed to be of the same class.
[…]
If $c_{q}$ matches the true class of the query instance, it is counted as a correct classification; otherwise, it is counted as an incorrect classification. In order to improve the quality of the similarity score, we perform min-max scaling on each image pair and shift them to a center grid to align them as best as possible. For the similarity computations in the classification tasks, the query is also shifted in eight different positions to test for the best alignment: up, down, left, right, and with 15 and 25 degrees clockwise and counterclockwise rotations. The hyperparameters of the circle radius, the density of pixels generated by the GMM, and the number of Gaussian components were optimized using grid-search based on classification accuracy across 100 iterations. We found the best performance with radius = 1.6, density = 300, and number of components = 10. Additionally, we set $\beta$ = 1.4.
Figure 2: Visual Turing tests of the output characters generated from our AGP-VAE pipeline. The set of characters drawn by our model is B for within alphabet, A for same class, and B for unconstrained.
5.2 Generative Tasks
The primary generative tasks in the Omniglot challenge are as follows:
  1. 1.
    Generating new exemplars of a particular class.
  2. 2.
    Generating new classes consistent with a particular alphabet given a starting set.
  3. 3.
    Generating entirely new classes (unconstrained).
For Task (1), only a single instance is used for the entire task. For Task (2), one instance per class from a given alphabet is allowed. Following (Lake et al., 2019), we use ten different instances. Similarly, Task (3) uses the same technique as Task (2) except that classes are randomly sampled across alphabets. Overall, the approach between these three tasks is similar and the only difference is in the starting instance(s) that are used.
There are three steps to generating new variations of exemplars from single instances. First, we synthetically increase the amount of training data by generating a larger number of AGPs per class. A key feature of this approach is that a range of prototypes with varied manifestations of abstraction can be generated by specifying a range of different number of components in the GMM for each class. This increases variation at the subparts level to diversify the training set. Second, we train a VAE across all the synthetic data to learn a continuous space of prototypes derived from the starting set. Lastly, we use a post-processing topological skeleton technique (Lee et al., 1994; Zhang, 1997) to refine the generated outputs. This layer denoises the reconstructed outputs of the VAE to ensure quality stroke images of the new class variants. The following sections describe each step in detail, and the pseudocode is shown below in Algorithm 1.
Input:
$\mathcal{X}\leftarrow$ set of single instances from $N$ classes
$K\leftarrow$ range of Gaussian Components
GMM, VAE $\leftarrow$ trainable models
Skel $\leftarrow$ topological skeletonization function
Output:
$\Phi\leftarrow$ final training set of prototypes
$z\leftarrow$ reconstructed latent variables from trained VAE
$v\leftarrow$ final generated variant
for $i\in N$ do
       for $k\in K$ do
             train a $GMM_{i,k}(x_{i})$ using $k$ components
$\mathcal{P}\leftarrow$ sample from $GMM_{i,k}$
$\Phi_{i}\leftarrow\Phi_{i}\cup\{\mathcal{P}\}$ append $\mathcal{P}$ to set of class prototypes
       end for
end for
$\Phi\leftarrow\Phi\cup\{\Phi_{i}\}$ append class train set to final train set
train VAE($\Phi$)
$z\leftarrow$ sample and reconstruct VAE latent variables
$v\leftarrow$ Skel($z$) postprocess reconstructed image
Algorithm 1 Generating New Variants
5.2.1 AGP Training Set
The approach from section 5.1.1 is used to generate more AGPs for each concept to synthetically increase the size of a training set $\Phi=\{\Phi_{1},\Phi_{2},\ldots,\Phi_{N}\}$. This training set consists of a larger set of AGPs where $\Phi_{i}=\{\mathcal{P}_{1},\mathcal{P}_{2},\ldots,\mathcal{P}_{D}\}$, containing $D$ new variants for each class, specified with a different value of $k$ components. An equal number of variants is generated for each value of $k$, so that there are $D/K$ variants for $k\in K$. In our pipeline, we generate $D$ = 500 AGPs per class with $K$ = {6,7,8,9 10} components. So for each class, we create 100 new instances for each k value.
5.2.2 VAE Interpolation
After generating the prototype training sets $\Phi_{i}$ for each class $i\in N$, the next step is to create continuous variations amongst these prototypes. To accomplish this, a VAE is trained across $\Phi=\{\Phi_{1},\Phi_{2},\ldots,\Phi_{N}\}$ which is the enumeration of all prototype training sets, to learn a latent space representation that captures the underlying structures of these abstracted prototypes. The latent variables $z$ are sampled accordingly to encourage semantic mixing between prototypes which are then decoded into the reconstructed variant images. For our generated outputs, we specify a convolutional VAE with the following details. The encoder of our model consists of two convolutional layers: a 32-filter 3x3 convolution followed by a 64-filter 3x3 convolution, both with a stride of 2 and ReLU activation. The decoder reshapes the latent vector to a 7x7x32 tensor, upsampled using two transposed convolutional layers with 64 and 32 filters respectively, both with 3x3 kernels, a stride of 2, and ReLU activation. The final layer is a transposed convolution with a single filter, a 3x3 kernel, and stride of 1, outputting the reconstructed image. For training, the model uses the Adam optimizer (Kingma & Ba, 2014) with a learning rate of 1e-4. The loss function is computed as a combination of binary cross-entropy and the KL divergence between the learned latent distribution and the prior distribution. Finally, we train the model using a batch size of 32 for 50 epochs.
5.2.3 Topological Skeleton Refinement
The final step in this pipeline is a post-processing technique based on the work of (Lee et al., 1994; Zhang, 1997) on topological skeletons. Skeletonization is used often in image processing and computer vision to reduce the thickness of binary objects to one-pixel-wide representations while preserving the topological properties of objects. This step refines the reconstructed output images generated by the VAE and emphasizes the stroke-like properties of Omniglot characters. After each reconstructed image from the VAE is skeletonized, the final result is a collection of generated variants of characters for each task.
6 Results
6.1 Classification Tasks
The classification accuracy of the proposed approach is evaluated based on the number of correct responses averaged over 1000 trials in both 5-way and 20-way one-shot tasks. We test this in the context of unconstrained classes independent of alphabets, as well as a more challenging within-alphabet classification task. We compared our one-shot classification approach using the Tversky-inspired similarity metric to a baseline euclidean distance metric calculated by the standard mean squared error (MSE). Given our simple binary image stimuli, we found MSE as the most relevant baseline over other metrics like cosine similarity, which evaluates similarity between high-dimensional embeddings and feature vectors, and structural similarity index (SSIM), which unnecessarily considers texture, luminance, and contrast not present in line-based binary strokes. Results are summarized in Table 1.
Table 1: Average Classification Scores
Euclidean Distance Ours
5-Way Unconstrained 34.5% 95.1%
5-Way Within 31.2% 86.6%
20-Way Unconstrained 19.3% 84.2%
20-Way Within 8.9% 71.0%
6.2 Generative Tasks
A ”visual Turing test”, as described in (Lake et al., 2013) is used to assess the quality of the generative outputs of the model. In this test, a set of characters produced by a human is displayed next to a set produced by the model. Human judges then try to identify which set was drawn by a human, and which set was generated by the model, see Figure 2 and Appendix A. Our generative approach is evaluated based on the identification accuracy of the 20 human judges recruited online. The ideal performance is 50 percent, indicating that the judges cannot distinguish between characters produced by the human and the model, and the worst-case performance is 100 percent. Ten question sets with four instances from the human and four instances from our model were created for each of the three tasks (total of 30 sets). Additionally, we asked follow up questions after displaying each set of images to probe whether the machine’s outputs could potentially surpass the quality of human generated characters. These questions were phrased as the following: (1) “Which set represents a better job of making four new examples of the given character?”, (2) “Which set represents a better job of making four new characters that fit the given alphabet?”, and (3) “Which set represents a better job of creating four new characters?”
Generating New Concepts from Type For the evaluation of generating new characters belonging to an alphabet, the identification accuracy across judges was (M = 52.00%, SD = 14.73%. Min = 40.00%, Max = 90%). The preference for the machine-made was (M = 49.00%, SD = 15.18%, Min = 20.00%, Max = 80.00%).
Generating New Exemplars For the set of images corresponding to generating new exemplars of a particular class, the average identification accuracy across judges was (M = 57.50%, SD = 14.82%. Min = 30.00%, Max = 80%). The preference for machine-made in this specific task was (M = 59.50%, SD = 12.76%, Min = 30.00%, Max = 80.00%).
Generating New Concepts (Unconstrained) For the final task of generating entirely new concepts independent of alphabet, the identification accuracy across judges was (M = 47.50%, SD = 12.01%. Min = 30.00%, Max = 70%). The preference for the machine-made in this task was (M = 57.50%, SD = 12.93%, Min = 30.00%, Max = 80.00%).
Overall Results for Generative Tasks The overall identification accuracy and preference scores averaged across all tasks and judges are shown in Table 2. In addition, Figure LABEL:fig:individual provides a breakdown to reveal the subjective evaluations of each individual judge. The primary measure of identification accuracy is quite close to chance performance meaning that judges were largely unable to tell which characters were made by human or machine. Accordingly, one can conclude that the model-generated characters were fully convincing. In terms of the preference measures, human-made characters were more frequently preferred for new exemplar generation and equally preferred for the remaining tasks. Notably, preference for machine-generated characters was overall slightly higher than human-drawn characters which could merit further exploration in terms of AI-generated content.
Table 2: Descriptive Statistics for Average Scores Across Judges
Identification Accuracy Preference
Mean 52.33% 55.33%
SD 8.17% 8.27%
Min 40.00% 43.33%
Max 63.33% 70.00%
7 Conclusion
We present a novel approach for addressing the Omniglot challenge using Abstracted Gaussian Prototypes in conjunction with a VAE for generative tasks. AGPs leverage GMMs to build representative prototypes for each concept by abstracting the subparts of the single available instance for each class. The AGP functions as a generative image segmentation method applied to cluster the pixels of an image which offering a simple yet powerful method for encoding higher level representations that capture a quasi-structural model of ’what and where’ for visual concepts from minimal data. One-shot classification is achieved by applying Tversky’s set-theoretic similarity metric to compare AGPs. A novel AGP-VAE pipeline employs VAEs to utilize AGPs with different component numbers to generate diverse and compelling (i.e., consistent, yet original) variants that judges found indistinguishable from human drawings.
While our approach presents promising contributions to one-shot learning, there are several limitations to be acknowledged. First, we acknowledge that these are not state-of-the art results for one-shot classification; however, we are not aware of another approach that performs this well under the strict interpretation of one-shot learning as being fully from scratch. The classification performance should also be considered in the context of the model’s excellent performance on the generative tasks in the challenge. In terms of potential wider impact of our approach, our method currently only handles line-based images with binary (on/off) pixels. The scalability to more complex natural images (i.e., ImageNet, CIFAR, etc.) would require new approaches to handle more features such as color. Furthermore, computational challenges from AGP generations and Tversky calculations can arise when working with non-binary images. Future research directions should aim to address these limitations and further refine the AGP framework for broader and more robust applicability in the field of one-shot learning.
In sum, we present AGPs as a novel approach that achieves high performance on one-shot classification task and proves impressively adaptability to the generative tasks of the Omniglot challenge. Critically, this degree and breadth of success is achieved without high model complexity, without slow and demanding computation, without lack of transparency, without the need for external pre-training, and without invoking a complex symbol system for explicit structural recoding. Ideal takeaways from this work include: (1) the value of the computational cognition framework for productive crosstalk between cognitive science and ML, (2) the potential for approaches that are intermediate in nature between the poles of the statistical and symbolic frameworks, (3) the future potential of the design principles of the abstracted Gaussian prototype and the AGP-VAE pipeline for ’truly’ one-shot learning and generative tasks without having to first learn how to learn.
Appendix A Appendix
All code, question sets, and data can be accessed at https://github.com/bosonphoton/AbstractedGaussianPrototypes. The full set of questions from our ”visual Turing tests” given to the human judges. The set of character outputs generated by our model are (from left to right, top to bottom):
Within Class: A A B A B B A B B A
Same Class: B A A A B A B B A A
Unconstrained: A B B A A B A B A A